# No-prior Bayes reIMagined: probabilistic approximations of possibilistic inferential models

### v2

@@ F.1 Bivariate normal correlation
F.2 Gamma shape and scale
Inference on the parameters of a gamma model is a classical problem. When the shape parameter is unknown, this model lacks the group transformation structure that makes specification of default priors having suitable reliability properties non-trivial. The go-to no-prior Bayes solution invokes Jeffreys prior on the shape–scale parameter pair. Alternatively, a possibilistic IM solution is readily available and doesn’t require any prior specification and has reliability guarantees. This IM solution is somewhat expensive to compute, however, at least if a naive approach is taken. Martin, (2025) showed how this naive approach can be drastically improved and it’s on this more sophisticated/efficient approach that my solution is based. Indeed, here I’m proposing the inner probabilistic approximation of that possibilistic IM, and the sampling from that “posterior distribution” is based on the strategy put forward in Martin, (2025).

### v3

@@ G.1 Bivariate normal correlation
G.2 Gamma shape and scale
Inference on the parameters of a gamma model is a classical problem. When the shape parameter is unknown, this model lacks the group transformation structure that makes specification of default priors having suitable reliability properties non-trivial. The go-to no-prior Bayes solution invokes Jeffreys prior on the shape–scale parameter pair. Alternatively, a possibilistic IM solution is readily available and doesn’t require any prior specification and has reliability guarantees. This IM solution is somewhat expensive to compute, however, at least if a naive approach is taken. Martin, 2025a showed how this naive approach can be drastically improved and it’s on this more sophisticated/efficient approach that my solution is based. Indeed, here I’m proposing the inner probabilistic approximation of that possibilistic IM, and the sampling from that “posterior distribution” is based on the strategy put forward in Martin, 2025a .